OpenStudy (beautifulpansophy):

A positive point charge q = +2.50 nC is located at x = 1.20 m and a negative charge of -2q = -5.00 nC is located at the origin. (a) Find a symbolic expression for the potential on the x – axis at an arbitrary point P between the two charges. (b) Find the electric potential at x = 0.600 m. (c) Find the point along the x –axis between the two charges where the electric potential is zero.

OpenStudy (michele_laino):

here we have to apply the superposition principle, namely the requested potential, is given by the algebraic sum of the single potentials:  \[\huge V\left( P \right) = K\frac{{ - 2q}}{x} + K\frac{q}{{d - x}}\] where, as usual, \(K=9 \cdot 10^9\)